OpenStudy (anonymous):

solve the polynomial equation 2x^4=16x^2

OpenStudy (anonymous):

x=0,+2*sqrt(2),-2*sqrt(2)

OpenStudy (anonymous):

Move the RHS to the LHS \[2x^4-16x^2=0\]The result can be factored.\[2 x^2 \left(x^2-8\right)=0\]